A thief is running away on a straight road on a jeep moving with a speed of 9m/s. A policeman chases him on a motor cycle moving at a speed 10m/s. if the instaneous separation of jeep from the motor cycle is 100m, how long will it take for the policeman to catch the theif?

Grade:12

2 Answers

Hari Sundar Pyda
6 years ago
Velocity of thief= 9m/s= Vt
Velocity of policeman= 10m/s= Vp
Initial distance between them= 100m
We know that relative velocity Vpt= Vp-Vt
= 10-9= 1m/s
Time taken= distance/speed 
= 100/1
=100 seconds 
= 1 minutes 40 seconds.
YOGITA GARBYAL
6 years ago
relative velocity of policeman with respect to theif;
VR=VP-VT
where
VP=velocity of police
Vt=velocity of theif
HANCE,
 Vr= 10-9=1 m\s
TIME TAKEN TO CATCH
t= distance of saperation\ relative  velocity
  100\1=100sec